Enorym:

Help!! \(2^{3x 1}=5^{5-x}\)

Enorym:

@umm @Elsa213 @ThisGirlPretty

ThisGirlPretty:

@563blackghost @Shadow

sillybilly123:

sure you got that right?!?! screengrab?

Enorym:

\( 2^{3x+1}=5^{5-x}\) I forgot the addition, sorry @sillybilly123

sillybilly123:

so take logs, natural logs if you like \(2^{3x+1}=5^{5-x}\) \(\ln ( 2^{3x+1} )= \ln ( 5^{5-x})\) \((3x+1) \ln 2 = (5-x) \ln 5\) And so on, solving for x

Enorym:

I got there, then I did \[(3x)ln2+ln2=(5)ln5+(-x)ln5\]

sillybilly123:

sure....and park the x's on the LHS and the non-x's on RHS

Enorym:

\[3x ~ln~2+x ln~5= 5ln~5-ln~2\]

sillybilly123:

\(3x \ln~2+x \ln~5= 5 \ln~5-\ln~2\) put a slash in front of the *ln* in tour Latex, smartens it up, like this \(x (3\ln~2+ \ln~5)= 5 \ln~5-\ln~2\) \(\implies x = \dfrac{5 \ln 5-\ln 2}{3\ln 2+ \ln 5} \)

sillybilly123:

the slash also creates a space afterwards so you don't need to use copious ~'s

Enorym:

Ah, thanks

sillybilly123:

mp/np

Enorym:

Too many abbreviations

sillybilly123:

you can do more stuff with it like, eg just taking the numerator: \(5 \ln 5-\ln 2 = \ln \dfrac{5^5}{2}\) but the next simplification is the calculator....turning x into a simple number

Enorym:

Ah, I am not allowed to use a calculator for this part

sillybilly123:

well i think that means you're done.

Enorym:

Thanks a lot, sillybilly

sillybilly123:

TU

Enorym:

A lot of abbreviations, I think I may need to update myself in those.